# PhishLang A Lightweight, Client-Side Phishing Detection Framework using MobileBERT for Real-Time, Explainable Threat Mitigation

### v2

@@ Abstract.
In this paper, we introduce PhishLang, an open-source, lightweight language model specifically designed for phishing website detection through contextual analysis of the website. Unlike traditional heuristic or machine learning models that rely on static features and struggle to adapt to new threats, and deep learning models that are computationally intensive, our model leverages MobileBERT, a fast and memory-efficient variant of the BERT architecture, to learn granular features characteristic of phishing attacks. PhishLang operates with minimal data preprocessing and offers performance comparable to leading deep learning anti-phishing tools, while being significantly faster and less resource-intensive. Over a 3.5-month testing period, PhishLang successfully identified 25,796 phishing URLs, many of which were undetected by popular antiphishing blocklists, thus demonstrating its potential to enhance current detection measures. Capitalizing on PhishLang’s resource efficiency, we release the first open-source fully client-side Chromium browser extension that provides inference locally without requiring to consult an online blocklist and can be run on low-end systems with no impact on inference times. Our implementation does not only outperform prevalent (server-side) phishing tools, but is significantly more effective than the limited commercial client-side measures available. Furthermore, we study how PhishLang can be integrated with GPT-3.5 Turbo to create ”Explainable blocklisting” — which upon detection of a website, provides users with detailed contextual information about the features that led to a website being marked as phishing.
@@ 1. Introduction
To address these issues, we introduce PhishLang, an efficient and transparent phishing detection framework utilizing MobileBERT, a lightweight language model. PhishLang employs advanced linguistic capabilities to identify subtle phishing patterns in the website source code, thus moving beyond static feature sets. Our model operates with significantly lower resource requirements—using up to almost seven times less memory than the next resource efficient model (DistilBERT), while outperforming several ML-based models in detecting new attacks, particularly those with evasive features. Unlike most ML-based phishing detection approaches, we also evaluate PhishLang against various realistic adversarial attacks and implement patches to enhance the model’s robustness. Additionally, we integrate our model with GPT-3.5 Turbo (OpenAI, 2023) to create “Explainable Blocklisting,” a feature that provides users with contextual explanations on why a phishing website was blocked. This approach not only helps users understand and engage with the protection mechanisms more effectively but also educates them on recognizing phishing attempts independently when such protections are unavailable or fail.
@@
Section 4 outlines the development and implementation of our real-time phishing detection system, PhishLang, which continuously monitors live URLs from Certstream to detect new phishing threats (Section 4.1). To benchmark our model’s effectiveness against existing measures, we also compare PhishLang’s performance with traditional deep learning-based phishing models (in Appendix), as well as (Section 9.1) established anti-phishing tools, with a focus on detecting evasive phishing tactics (Section 4.2). We then evaluate the robustness of our model against adversarial phishing websites by testing sixteen realistic attack scenarios that manipulate website source code, developing patches as countermeasures (Section 5). In Section 6, we introduce a novel “Explainable Blocklist” mechanism that pairs PhishLang with GPT-3.5 Turbo to generate phishing alerts that include contextual explanations, which can help users understand why a website was flagged. This approach not only enhances the credibility of phishing warnings but can also deepen users’ understanding of various phishing techniques. Finally, in Section 7, we highlight our client-side phishing detection extension. The primary contributions of our work are:
@@
  4. (4)
    We build a system, “Explainable blocklisting,” which utilizes GPT 3.5T on PhishLang’s predictions to provide users with context-aware information on the features that made a website be detected as phishing.
@@ 3.2.1. Groundtruth
To establish a ground-truth dataset for training our classifier, we utilized Lin at el’s PhishPedia dataset (Lin et al., 2021), which includes 30K phishing and 30K benign websites. PhishPedia is not only the largest publicly available phishing dataset but is also notable for its reliability, having been used in several prior literature (Liu et al., 2022; Guo et al., 2021). Each entry in this dataset typically includes associated metadata, such as HTML source code, screenshots, and OCR text. The phishing URLs were originally obtained from OpenPhish (Openphish, [n. d.]), a widely recognized antiphishing blocklist. These URLs underwent preprocessing to eliminate inactive sites and false positives, and the researchers also implemented strategies to counteract evasive tactics like cloaking and conducted manual verifications to correct any inaccuracies in target brand representations. Conversely, the benign URLs were sourced from the Alexa Rankings list (ale, 2020), a (now defunct) online database that ranked websites based on popularity. For the purposes of training our model, we only required the HTML source file, such that it can be parsed into a collection of relevant segments as previously detailed in Section 3.1. Out of the 30K phishing entries, 22,419 had an HTML file, whereas the same number for benign URLs was 26K. We further analyzed this dataset beyond targeted organizations (Lin et al., 2021), to obtain the distribution of HTML tags, libraries and frameworks used, and JavaScript APIs and function calls. Table 1 illustrates the distribution of HTML tags in both our phishing and benign websites. Notably, phishing websites contain fewer tags than benign websites. This is because phishing websites primarily focus on capturing user credentials, whereas benign websites need to offer a wide range of functionalities, requiring a more extensive and complex HTML structure. We also used Wappalyzer (Chorsley, 2024) and Esprima (esp, [n. d.]) to identify the frameworks and JS functions used by and present respectively in these websites. For brevity, we include this analysis in Section 9.2 of the Appendix. Finally, we are interested in identifying whether PhishLang can detect evasive phishing attacks, although our model cannot categorize the types of evasion directly. Therefore, we applied separate heuristics, derived from specific studies (Oest et al., 2020a; Team, 2022; Auth0, 2021; Liang et al., 2016; Ventures, 2019), to classify the phishing URLs in the Phishpedia dataset into five categories: i) Regular phishing attacks (18,401), ii) Behavioral JS Evasions (715), iii) Clickjacking attacks (1,726), iv) Exploiting DOM (536), and v) Text encoding attacks (1,041). Behavioral JS Evasions, as highlighted by Oest et al.(Oest et al., 2020a), include JS-based redirection triggered by user interactions, mouse movement detections, and popup interactions, exploiting weaknesses in several prominent blocklists. In contrast, DOM Manipulations involve altering the HTML Document Object Model (DOM) to evade detection, such as dynamically generating phishing content, hiding malicious elements with CSS, and modifying the DOM structure post-page load. Liang et al. (Liang et al., 2016) identified features of these tactics that bypass Google’s phishing page filter (GPPF), which are also transferable to other detection models. Text encoding attacks include obfuscating the character encoding, nested encoding, encoding to obfuscate JavaScript (JS), and white space obfuscation, hiding malicious code within seemingly harmless text by manipulating how the text is interpreted by different browsers. We use 70% of our ground truth data for training the model and reserve 30% for testing. The model is trained using 5-fold cross-validation, with each fold undergoing 10 epochs of training.
@@ 3.2.2. Choosing the optimum model
The performance of these models is illustrated in Table 2. We found Llama 2 to have the best performance out of all the models tested, with an F1 score of 0.96. However, its median inference time of  34s is impractical for real-world usage where a phishing detection tool would need to process hundreds of thousands of URLs every day, and it has the highest median memory usage at 4,873.47MB. While the other two LLMs, Bloom and T5, had lower median inference times of 7.49 and 12.40 seconds respectively, they also performed worse (F1 scores of 0.88 and 0.74, respectively) and had high memory requirements (7351.89MB for Bloom and 1,279.49MB for T5). GPT-2 had the lowest median inference time (1.81 seconds) but also had the worst performance among the LLMs (F1 0.68) and required 922.41MB of memory. Moreover, all four LLMs required inference over GPUs (while they can be used for inference using the CPU, the prediction times will be much slower), and consequently required a large amount of VRAM for said inference per sample. Requiring GPU to provide good inference time is not ideal in a practical setting, as most consumer-end devices where the inference needs to occur (such as portable/mobile systems) might not have GPUs.
@@
When considering the Small Language Models (SLMs), MobileBERT had the highest accuracy level at 0.96, with both precision and recall scoring 0.95 and 0.96, respectively. It also demonstrated the fastest prediction times among the SLMs, with a median of 0.39 seconds per prediction, and required the least memory usage at 74MB among all models tested. These attributes make MobileBERT highly suitable for real-time phishing detection, balancing high performance with low computational resource requirements, which is ideal for deployment on devices with limited hardware capabilities, such as mobile and portable systems. Its efficient use of DRAM instead of requiring a GPU makes it even more attractive for widespread, scalable deployment. In comparison, DistilBERT achieved an accuracy of 0.94, with both precision and recall scoring 0.94. While not the quickest, its prediction times were reasonably efficient, with a median of 0.85 seconds per prediction and significantly higher memory usage of 502MB compared to MobileBERT. DeBERTa-base, another SLM, delivered lower precision and recall, also fell short in time efficiency, and had a higher memory usage of 1,341MB. FastText, though more compact with a memory usage of 201MB, suffered from reduced accuracy (0.62) and had a median prediction time of 1.43 seconds. Meanwhile, TinyBERT, the smallest model tested, offered quicker results at 0.78 seconds but at the cost of lower performance metrics, requiring 495MB of memory. All SLMs could provide inference over CPUs, making them suitable for deployment on various platforms. Also, due to the unavailability of commercial LLMs such as ChatGPT and Claude as local implementations, fine-tuning and testing on these models would incur significant costs for training and, especially for inference, given the millions of URLs an anti-phishing tool needs to process daily. However, we compare ChatGPT (3.5 Turbo) and GPT 4 with our model, and other ML-based phishing detection tools in Section 8 in the Appendix. Thus, considering all the tested language models, based on a combination of high precision and recall, fast prediction times, and efficient memory usage, MobileBERT was the most suitable choice for our framework and was thus chosen as the final language model for our PhishLang framework.
@@ 5.3. Advesarial impact on phishing categories
6. Explainable blocklisting
Figure 2. Example of PhishLang’s explainable blocklisting warning page showcasing contextual details about suspicious features present in the detected Facebook phishing page. Note that the website URL is not redacted since it is already inactive.
Figure 3. Framework for generating explainable features
Security systems often rely on warning dialogs to convey critical information, particularly when blocking malicious content or changes in system configurations that could impact usability (Raja et al., 2011). However, frequent and generic warnings offer little insight into actual threats, leading to user desensitization and ignored alerts (Voronkov et al., 2017; Akhawe and Felt, 2013; Krol et al., 2012). This problem is prevalent in antiphishing systems integrated into web browsers (saf, 2020; mca, 2022), which typically indicate a dangerous website without explaining the specific reasons for flagging it. Evidence suggests that providing contextual information about threats can significantly enhance user understanding and trust (Desolda et al., 2023; Nguyen et al., 2021), yet this approach is rarely adopted in commercial antiphishing tools, leaving users uninformed. Our goal is to leverage PhishLang to develop a system that not only identifies suspicious features in phishing website detection but also provides contextual explanations for why these features are considered suspicious. While Desolda et al. (Desolda et al., 2023) have explored how explainable blocklisting can aid users in recognizing phishing websites, their system is limited to analyzing a narrow set of URL features. Given the emergence of new evasive techniques, these features alone may not effectively indicate phishing attacks. Additionally, their approach relies on generating static messages based on these features, offering users limited context. Recognizing the shortcomings of current warning systems, our method utilizes the capabilities of GPT-3.5 Turbo in conjunction with our MobileBERT-trained model to deliver comprehensive contextual information when users encounter phishing alerts. Instead of simply notifying users that a website is malicious, our approach analyzes the specific suspicious features of the flagged site and highlights their locations, providing users with a deeper understanding of the threat.
6.1. Generating explainable features
When our MobileBERT model detects a website as phishing, it utilizes LIME to pinpoint the most influential features in its prediction, in an approach similar to the one illustrated in Section 6.3. We consider identifying the top three tags that most significantly influenced the model’s decision and select the tags within the website’s parsed source code that contained those phrases. We chose three tags since we wanted the model to generate explanations for at least three features. However, this value can be increased or decreased as required. These tags, together with the entire parsed source code and the site’s URL, are integrated into a custom prompt that we provide to GPT-3.5T using the OpenAI API. Our framework is illustrated in Figure 3. The purpose of this prompt is to enable the language model to create a dialogue that elaborates on the contextual significance of these suspicious features and their exact locations on the website. Figure 2 illustrates a phishing page designed to imitate Facebook’s Privacy Center that was detected by PhishLang, and the subsequent detailed warning was generated. The system identifies several key indicators of phishing on the page. Firstly, it notes the use of language designed to alarm/rush the user, a common tactic in phishing schemes aimed at prompting users to disclose information impulsively (Alsharnouby et al., 2015). Secondly, the tool highlights that the links on the sidebar are non-functional, another tactic that is used by attackers to make the website look legitimate (Li et al., 2019). Thirdly, PhishLang flags noticeable grammatical mistakes which are not typical of professional communication from a well-established company like Facebook. For each feature, the system pinpoints where it appears on the webpage and includes a screenshot, enabling users to visually match the suspicious elements with the explanations provided.
Figure 4. Prompt used for generating explainable warnings
6.2. Choice of generative model
For automated explainable blocklisting, it is crucial that the AI model can generate coherent and contextually appropriate human dialogue. Our decision to use GPT 3.5T, a powerful open-domain chatbot, was informed by several limitations observed in locally deployable LLMs. Firstly, while the MobileBERT architecture is adequate for binary classification tasks, its capabilities in text generation (along with its parent model - BERT) are insufficient (Chan and Fan, 2019). More robust models such as LLaMA 2 and T5, although capable of producing meaningful explanations, were deemed impractical for real-world application due to their high resource consumption and slow performance. For instance, the smallest LLaMA 2 model, with 7 billion parameters (TheBloke, [n. d.]), required an average of 72 seconds to generate an explanation on our test system. In contrast, GPT-3.5T not only outperforms these local LLMs in both speed and output quality—taking an average of just 4 seconds to generate an explanation—but it also presents a more cost-effective solution. The operational cost of GPT-3.5T is significantly lower than that of the GPT-4 model, averaging about \$1 for 2,000 explanations.
6.3. Reliability of generated content
Large Language Models such as GPT 3.5T inherently produce non-deterministic content, meaning the output can vary significantly even when provided similar input conditions across different instances (Kocoń et al., 2023). Furthermore, these models are susceptible to ”hallucinations” — instances where they generate information that, while plausible, is either incorrect or unverifiable (Alkaissi and McFarlane, 2023). Given these characteristics, it is imperative to conduct thorough pre-processing and verification of the generated explainable warnings to ensure the accuracy and reliability of the model’s capability to identify and describe suspicious features accurately.
Pre-processing content: To ensure consistency in GPT-3.5T’s output, we prompted it to compile responses as a structured JSON with key-value pairs for the feature name, location, and description (i.e., the context of the suspicious feature within the website). This data is parsed by a driver to generate a warning displayed to the user. We utilized our codebook (https://github.com/UTA-SPRLab/phishlang) to define specific feature terms, instructing GPT-3.5T to select from this predefined list, ensuring the model does not invent alternate terms for the same feature. Only visible website features are considered, omitting features like ”Session Hijacking attempt,” ”IP Tracking,” or ”JavaScript redirection.” If GPT-3.5T identifies features not in our list but significantly supporting detection, we encouraged their identification. During the evaluation, GPT-3.5T identified three new features: ”Grammatical errors” (mistakes in spelling, grammar, and punctuation), ”Poor security indicators” (e.g., use of HTTP instead of HTTPS), and ”Mismatched URLs” (URLs not matching the official URL of the targeted organization). These were subsequently added to our feature list.
Human review: LLMs can hallucinate information (Alkaissi and McFarlane, 2023). Despite our structured prompt, we needed to evaluate the reliability of ChatGPT’s output. Two coders manually reviewed a sample of explainable warnings generated for 100 randomly selected websites from our training dataset. Coders were provided with a screenshot of the website, its parsed source code, and its URL. They assessed: a) if GPT-3.5T’s features were present on the website, b) if the feature location was correct, c) if the feature description was accurate, and d) if any additional features were missed. The coders achieved a Cohen’s Kappa inter-reliability agreement of 0.78. After reviewing the websites, they resolved disagreements through discussion. In total, 401 feature points were identified across 100 websites. The targeted organization imitated by the phishing website was correctly identified in all cases. Coders found 383 ( 96%) features correctly identified, with the correct location for 339 ( 84%) of them. Among the 18 incorrect identifications were non-sensitive text fields (9), hallucinated two-factor authentication modules (5), and multiple brand texts (4).

### v3

# PhishLang: A Real-Time, Fully Client-Side Phishing Detection Framework Using MobileBERT

@@ Abstract.
In this paper, we introduce PhishLang, the first fully client-side anti-phishing framework built on a lightweight ensemble framework that utilizes advanced language models to analyze the contextual features of a website’s source code and URL. Unlike traditional heuristic or machine learning approaches that rely on static features and struggle to adapt to evolving threats, or deep learning models that are computationally intensive - our approach utilizes MobileBERT, a fast and memory-efficient variant of the BERT architecture, to capture nuanced features indicative of phishing attacks. To further enhance detection accuracy, PhishLang employs a multi-modal ensemble approach, combining both the URL and Source detection models. This architecture ensures robustness by allowing one model to compensate for scenarios where the other may fail, or if both models provide ambiguous inferences. As a result, PhishLang excels at detecting both regular and evasive phishing threats, including zero-day attacks, outperforming popular anti-phishing tools, while operating without relying on external blocklists and safeguards user privacy by ensuring that browser history remains entirely local and unshared. We release PhishLang as a Chromium browser extension and also open-source the framework to aid the research community.
@@ 1. Introduction
To address these issues, we introduce PhishLang, an efficient and transparent phishing detection framework utilizing MobileBERT, a lightweight language model. PhishLang employs advanced linguistic capabilities to identify subtle phishing patterns in the website source code, thus moving beyond static feature sets. Our model operates with significantly lower resource requirements—using up to almost seven times less memory than the next resource efficient model (DistilBERT), while outperforming several ML-based models in detecting new attacks, particularly those with evasive features. Unlike most ML-based phishing detection approaches, we also evaluate PhishLang against various realistic adversarial attacks and implement patches to enhance the model’s robustness.
@@
Section 4 outlines the development and implementation of our real-time phishing detection system, PhishLang, which continuously monitors live URLs from Certstream to detect new phishing threats (Section 4.1). To benchmark our model’s effectiveness against existing measures, we also compare PhishLang’s performance with traditional deep learning-based phishing models, as well as (Section 3.3) established anti-phishing tools, with a focus on detecting evasive phishing tactics (Section 4.2). We then evaluate the robustness of our model against adversarial phishing websites by testing sixteen realistic attack scenarios that manipulate website source code, developing patches as countermeasures (Section 5). Finally, in Section 6, we highlight our client-side phishing detection extension. The primary contributions of our work are:
@@ 3.2.1. Groundtruth
To establish a ground-truth dataset for training our classifier, we utilized Lin at el’s PhishPedia dataset (Lin et al., 2021), which includes 30K phishing and 30K benign websites. PhishPedia is not only the largest publicly available phishing dataset but is also notable for its reliability, having been used in several prior literature (Liu et al., 2022; Guo et al., 2021). Each entry in this dataset typically includes associated metadata, such as HTML source code, screenshots, and OCR text. The phishing URLs were originally obtained from OpenPhish (Openphish, [n. d.]), a widely recognized antiphishing blocklist. These URLs underwent preprocessing to eliminate inactive sites and false positives, and the researchers also implemented strategies to counteract evasive tactics like cloaking and conducted manual verifications to correct any inaccuracies in target brand representations. Conversely, the benign URLs were sourced from the Alexa Rankings list (ale, 2020), a (now defunct) online database that ranked websites based on popularity. For the purposes of training our model, we only required the HTML source file, such that it can be parsed into a collection of relevant segments as previously detailed in Section 3.1. Out of the 30K phishing entries, 22,419 had an HTML file, whereas the same number for benign URLs was 26K. We further analyzed this dataset beyond targeted organizations (Lin et al., 2021), to obtain the distribution of HTML tags, libraries and frameworks used, and JavaScript APIs and function calls. Table 1 illustrates the distribution of HTML tags in both our phishing and benign websites. Notably, phishing websites contain fewer tags than benign websites. This is because phishing websites primarily focus on capturing user credentials, whereas benign websites need to offer a wide range of functionalities, requiring a more extensive and complex HTML structure. We also used Wappalyzer (Chorsley, 2024) and Esprima (esp, [n. d.]) to identify the frameworks and JS functions used by and present respectively in these websites. Finally, we are interested in identifying whether PhishLang can detect evasive phishing attacks, although our model cannot categorize the types of evasion directly. Therefore, we applied separate heuristics, derived from specific studies (Oest et al., 2020a; Team, 2022; Auth0, 2021; Liang et al., 2016; Ventures, 2019), to classify the phishing URLs in the Phishpedia dataset into five categories: i) Regular phishing attacks (18,401), ii) Behavioral JS Evasions (715), iii) Clickjacking attacks (1,726), iv) Exploiting DOM (536), and v) Text encoding attacks (1,041). Behavioral JS Evasions, as highlighted by Oest et al.(Oest et al., 2020a), include JS-based redirection triggered by user interactions, mouse movement detections, and popup interactions, exploiting weaknesses in several prominent blocklists. In contrast, DOM Manipulations involve altering the HTML Document Object Model (DOM) to evade detection, such as dynamically generating phishing content, hiding malicious elements with CSS, and modifying the DOM structure post-page load. Liang et al. (Liang et al., 2016) identified features of these tactics that bypass Google’s phishing page filter (GPPF), which are also transferable to other detection models. Text encoding attacks include obfuscating the character encoding, nested encoding, encoding to obfuscate JavaScript (JS), and white space obfuscation, hiding malicious code within seemingly harmless text by manipulating how the text is interpreted by different browsers. We use 70% of our ground truth data for training the model and reserve 30% for testing. The model is trained using 5-fold cross-validation, with each fold undergoing 10 epochs of training.
@@ 3.2.2. Choosing the optimum model
The performance of these models is illustrated in Table 2.
Figure 2. Distritbution of: Libraries/Frameworks found in: A) Phishing websites, B) Benign Websites, and JS Main function calls found in: C) Phishing websites.
@@
We found Llama 2 to have the best performance out of all the models tested, with an F1 score of 0.96. However, its median inference time of  34s is impractical for real-world usage where a phishing detection tool would need to process hundreds of thousands of URLs every day, and it has the highest median memory usage at 4,873.47MB. While the other two LLMs, Bloom and T5, had lower median inference times of 7.49 and 12.40 seconds respectively, they also performed worse (F1 scores of 0.88 and 0.74, respectively) and had high memory requirements (7351.89MB for Bloom and 1,279.49MB for T5). GPT-2 had the lowest median inference time (1.81 seconds) but also had the worst performance among the LLMs (F1 0.68) and required 922.41MB of memory. Moreover, all four LLMs required inference over GPUs (while they can be used for inference using the CPU, the prediction times will be much slower), and consequently required a large amount of VRAM for said inference per sample. Requiring GPU to provide good inference time is not ideal in a practical setting, as most consumer-end devices where the inference needs to occur (such as portable/mobile systems) might not have GPUs.
@@
When considering the Small Language Models (SLMs), MobileBERT had the highest accuracy level at 0.96, with both precision and recall scoring 0.95 and 0.96, respectively. It also demonstrated the fastest prediction times among the SLMs, with a median of 0.39 seconds per prediction, and required the least memory usage at 74MB among all models tested. These attributes make MobileBERT highly suitable for real-time phishing detection, balancing high performance with low computational resource requirements, which is ideal for deployment on devices with limited hardware capabilities, such as mobile and portable systems. Its efficient use of DRAM instead of requiring a GPU makes it even more attractive for widespread, scalable deployment. In comparison, DistilBERT achieved an accuracy of 0.94, with both precision and recall scoring 0.94. While not the quickest, its prediction times were reasonably efficient, with a median of 0.85 seconds per prediction and significantly higher memory usage of 502MB compared to MobileBERT. DeBERTa-base, another SLM, delivered lower precision and recall, also fell short in time efficiency, and had a higher memory usage of 1,341MB. FastText, though more compact with a memory usage of 201MB, suffered from reduced accuracy (0.62) and had a median prediction time of 1.43 seconds. Meanwhile, TinyBERT, the smallest model tested, offered quicker results at 0.78 seconds but at the cost of lower performance metrics, requiring 495MB of memory. All SLMs could provide inference over CPUs, making them suitable for deployment on various platforms. Also, due to the unavailability of commercial LLMs such as ChatGPT and Claude as local implementations, fine-tuning and testing on these models would incur significant costs for training and, especially for inference, given the millions of URLs an anti-phishing tool needs to process daily. However, we compare ChatGPT (3.5 Turbo) and GPT 4 with our model, and other ML-based phishing detection tools in Section 3. Thus, considering all the tested language models, based on a combination of high precision and recall, fast prediction times, and efficient memory usage, MobileBERT was the most suitable choice for our framework and was thus chosen as the final language model for our PhishLang framework.
3.3. Comparison with other ML models
We compared the performance of PhishLang with other popular open-source ML-based phishing detection models. Two of these models rely on the visual features of the website: VisualPhishNet (Abdelnabi et al., 2020) and PhishIntention (Liu et al., 2022), one relies on both the URL string and HTML representation of the website: StackModel (Li et al., 2019), and one that relies on the semantic representation of the URL string only: URLNet (Le et al., 2018). Each of these models was tested on the 5K samples, which we had labeled earlier. We also include the commercially available models of ChatGPT (GPT 3.5T and GPT 4). Since the focus of PhishLang is not only on prediction performance but also on the overhead of the models, which is crucial for practical implementation, we add two metrics - Median Artifact size, i.e., the amount of data (source code, URL, image, etc. based on the model) that needs to be provided to the model for prediction, and Median Prediction Time, which is the median time required for the respective model to provide the prediction label. Also, since GPT 3.5T and 4 are general LLM models that can predict any text content, we pass both the full HTML and the parsed versions as artifacts for evaluation. Table 3 illustrates the performance of each of the models. We note that PhishIntention, despite showing the best performance among all the models, requires a relatively large artifact size of 348KB and has a longer prediction time of 10.7 seconds. PhishLang, on the other hand, shows strong performance metrics with the advantage of a significantly smaller artifact size (7KB) and faster prediction time (0.9 seconds), indicating a balance between high performance and operational efficiency. Visual PhishNet and StackModel show moderate performance in all metrics, with the former requiring a much larger artifact size and longer prediction time compared to StackModel. Finally, URLNet has the lowest performance metrics but benefits from the smallest artifact size (0.2KB) and shortest prediction time (0.7 seconds). In the case of the ChatGPT models, we found that GPT 3.5 was only able to process 822 phishing + 1031 benign samples when the full HTML was given, while GPT 4 was able to process 1,469 phishing + 2,147 benign samples due to the token size limit. We find that both models perform very well against the threats and are almost on par with both PhishLang and PhishIntention. Note that these GPT models are general models and were not specifically fine-tuned using phishing training data. Thus, it is possible that it can outperform our model if fine-tuned on appropriate data. However, restricting GPT models for commercial use and the overall expense required for training them could make them unfeasible for practical implementation.
Table 3. Comparison with other ML-based detection methods
Model Accuracy Precision Recall F1 Score Size Time
PhishLang 0.94 0.93 0.95 0.94 7KB 0.9
PhishIntention 0.96 0.94 0.96 0.95 348KB 10.7
Visual PhishNet 0.85 0.83 0.86 0.85 217KB 4.9
URLNet 0.73 0.72 0.74 0.73 0.2KB 0.7
StackModel 0.83 0.85 0.81 0.82 1KB 1.9
GPT 3.5T (HTML)* 0.87 0.88 0.86 0.87 42KB 2.1
GPT 4 (HTML)* 0.92 0.93 0.91 0.92 42KB 2.8
GPT 3.5T (Parsed) 0.90 0.90 0.88 0.88 7KB 1.6
GPT 4 (Parsed) 0.94 0.93 0.93 0.93 7KB 1.4
@@ 5.3. Advesarial impact on phishing categories
5.4. Evaluating misdetections
Figure 3. Example of an Instagram attack where the seemingly benign landing page (which was missed by PhishLang) led to a phishing page.
In this section, we evaluate the misdetections (False Positives and False Negatives) of PhishLang for both the vanilla model and after adversarial retraining on the 2.5K websites detected as phishing by PhishLang. We also analyzed the reasons for false positives and false negatives, observed whether adversarial improvements introduced additional false negatives, and monitored changes in the false positive and false negative rates for evasive attack categories post-adversarial training.
5.4.1. False Negatives
Out of 261 undetected websites by PhishLang, 82 had CAPTCHAs and 36 had QR codes, both common phishing evasion tactics (Vidas et al., 2013; Kang and Xiang, 2010). PhishLang successfully detected all phishing websites when QR code URLs were scanned and most CAPTCHA-protected sites, except two inactive ones. For QR codes, we used the pyzbar(Hudson, [n. d.]) library to automate scanning. However, solving CAPTCHAs programmatically can violate many websites’ terms of service (Sivakorn et al., 2016), so PhishLang detects phishing post-CAPTCHA. We also found 25 sites that mimicked legitimate brands but initially only had input fields without malicious behavior. PhishLang’s confidence was low on these landing pages, resulting in false negatives, but it detected phishing on the subsequent pages. Figure 3 shows an Instagram phishing page where PhishLang’s confidence increased from 0.41 on the first page to 0.98 on the second. LIME analysis revealed PhishLang was influenced by brand-related text and form fields but lacked other phishing indicators initially. To address this, our framework now inputs placeholder data on benign-marked sites, submits it, and reassesses the destination, which helped detect all previously missed sites. Additionally, 118 undetected websites were in non-English languages: Japanese (57), Spanish (36), German (18), and Italian (7). Figure 4 shows an example in Japanese. Using langdetect, we found our training data was predominantly in English (n=18,683) with minimal coverage of these languages (Japanese=500, German=660, Spanish=403, and Italian=85). This lack of diversity likely caused misclassifications. While most phishing websites are in English, non-English phishing scams are emerging (Darktrace, 2024).
Figure 4. Example of a non-English (Japanese) false negative sample
Post-Adversarial Improvements: Implementing the patches from Section 5 significantly improved performance. Only 2 out of 36 QR-code attacks, 5 out of 25 multi-page attacks, and 7 true-positive detections were misclassified as benign, despite a median confidence score of 0.44.
5.4.2. False Positives
Upon evaluating the 112 false positives, we identified 42 malformed websites with poorly structured text and input fields soliciting sensitive information. For example, Fig  7 shows a website with invalid characters and poorly designed input fields for personal information like name, email, and phone number, lacking clear instructions or context. These issues likely caused the site to resemble phishing samples, leading to incorrect classification by PhishLang. Additionally, 23 websites featured multiple empty links and malformed meta tags with several redirections. Figure 5 illustrates an example where poorly rendered input fields and empty links indicated an incomplete population of the website. Addressing these false positives without compromising PhishLang’s detection features is challenging. We also discovered 14 websites with simple login fields that did not transmit data and contained no other information. These sites were classified as false positives because they neither imitated any organization nor had a deceptive motive, despite fitting traditional phishing profiles. Our crawler detects websites as soon as they appear on Certstream, so attackers could later add more details to these sites.
Figure 5. An example of a false positive with poor layout
We found 21 websites with domain parking pages, likely detected due to their lack of substantial content, presence of multiple advertising links, and frequent redirections. Figure 6 shows one such example. To assess PhishLang’s robustness against parked pages, we used keywords typical of parked pages to select 1,000 parked websites from our total sample of 42.7 million domains. PhishLang flagged only 5 of them as phishing (median confidence of 0.27).
Figure 6. An example of a parked domain page that was flagged as a false positive
Lastly, we noted 12 websites lacking noticeable phishing artifacts, with low confidence scores (Median=0.52). The small number of such samples might not significantly impact the training outcome if included in the dataset. We plan to retrain the model after gathering enough samples matching this criterion.
Post-Adversarial Improvements: Implementing adversarial improvements discussed in Section 5, PhishLang correctly identified 28 of the websites with poor structural design as benign. Out of the 2.5k manually observed true positives, no new samples were changed from phishing to benign, increasing the true positive rate of our model (on the initial dataset) to 96.6%.
Figure 7. Example of a poorly designed (benign)website that was detected as a false positive